### v2

# Tractable Probabilistic Models for Investment Planning

Abstract
Investment planning in power utilities, such as generation and transmission expansion, requires decade-long forecasts under profound uncertainty. Forecasting of energy mix and energy use decades ahead is nontrivial. Classical approaches focus on generating a finite number of scenarios (modeled as a mixture of Diracs in statistical theory terms), which limits insight into scenario-specific volatility and hinders robust decision-making. We propose an alternative using tractable probabilistic models (TPMs), particularly sum-product networks (SPNs). These models enable exact, scalable inference of key quantities such as scenario likelihoods, marginals, and conditional probabilities, supporting robust scenario expansion and risk assessment.
This framework enables direct embedding of chance-constrained optimization into investment planning, enforcing safety or reliability with prescribed confidence levels. TPMs allow both scenario analysis and volatility quantification by compactly representing high-dimensional uncertainties. We demonstrate the effectiveness of the approach through a representative power system planning case study, illustrating its computational and reliability advantages over traditional scenario-based models.
{IEEEkeywords}
Generation expansion planning, Mixture models, Power system investment, Sum-product networks, Transmission expansion planning
1 Introduction
footnotetext: †  These authors contributed equally.
With the increasing and more tangible effects of climate change, policymakers are intensifying demands for cleaner power systems, compelling grid operators to accelerate the integration of renewable energy into their networks. While renewables provide a feasible pathway to decarbonization and meeting greenhouse gas reduction targets, their weather-dependent and stochastic nature introduces profound operational challenges. Reliability is paramount: supply interruptions can propagate through the economy, disrupt critical infrastructure, and undermine social welfare. At the same time, the variability of renewable sources increases the need for additional reserves, flexibility options, and grid reinforcements, all of which carry high costs that grow with penetration levels [edenhofer2011renewable, impram2020challenges]. To maintain cost-effectiveness and resilience, system operators must ensure electricity remains affordable while safeguarding against blackouts and extreme events. These dual requirements have made rigorous planning indispensable, with infrastructure investments carefully evaluated to minimize the risk of decisions that could compromise long-term security and resilience of supply.
Planning in power systems has long focused on determining the amount of new capacity to build, when to deploy it, and which technologies are most suitable to meet projected demand growth. Traditionally, modelling these generation and transmission expansion problems relies on deterministic assumptions, where long-term forecasts of demand, fuel prices, and technology costs are fixed inputs [taylor2015convex, kirschen2018fundamentals]. Over time, however, it became evident that such assumptions overlook the fundamental uncertainty associated with economic growth, policy developments, the impacts of climate change, and demand response [paragElectricityMarketDesign2016a]. Consequently, more recent approaches emphasize the direct incorporation of uncertainty into planning frameworks to ensure that expansion strategies remain viable under a wide range of possible futures.
To address these challenges, planning methodologies have evolved towards stochastic programming and robust optimization [ben2009robust, conejo2010decision, pflug2014multistage, hadjidimitriou2021mathematical], which provide systematic ways of evaluating investment decisions under uncertain demand growth and variable renewable generation [zakaria2020uncertainty, 10931726]. In parallel, structural changes in the sector are shifting attention from large-scale, centralized capacity additions toward the integration of distributed energy resources (DERs), such as rooftop solar, storage systems, and demand response, which require more decentralized and flexible planning approaches [wang2013stochastic, asensio2016joint, MOTTA2024401]. Together, these developments reflect a broader shift in power system planning, where the traditional focus on least-cost expansion is now complemented by the need to ensure system resilience and long-term reliability in the face of uncertain and rapidly changing operating conditions.
Tractable probabilistic models (TPMs) [zhang2021probabilistic, poon2011sum, kisa2014probabilistic, darwiche2002logical] are a family of models that represent complex, high-dimensional probability distributions in a compact form. Their distinctive property is that they support efficient and exact inference in polynomial time, allowing the computation of likelihoods, marginals, and conditional probabilities [choi2020probabilistic]. Contrasting with classical probabilistic models, where inference often relies on sampling or approximate methods, which may be computationally prohibitive in large-scale applications. While TPMs share expressive power with neural networks and mixture models, they avoid the intractability of inference accompanying such approaches.
Recent surveys and reviews highlight the need for tractable probabilistic modeling in complex, high-dimensional systems characterized by deep uncertainty. Such systems, including modern energy networks, require scalable, exact reasoning to manage the volatility introduced by renewable integration, policy shifts, and dynamic demand [lavenezianaCriticalReviewEnergy2023].
In this work, we address the challenges of long-horizon investment planning in power utilities, such as generation and transmission expansion planning, where operators must make decisions decades in advance under profound uncertainty. Forecasting the future energy mix and consumption patterns over such horizons is inherently complex, and current practice often relies on a limited set of discrete scenarios that represent key trajectories of exogenous variables. Statistically, this corresponds to a mixture of Dirac measures, which allows planners to capture a handful of plausible futures but offers little insight into the variability and volatility within each scenario. We propose replacing this paradigm with inference over TPMs integrated into the optimization problem. By compactly representing distributions over high-dimensional uncertainties, TPMs enable reasoning about both scenarios and the associated volatility, and the inference is representable in mixed-integer optimization. We illustrate this approach with a representative case study, demonstrating its feasibility for expansion planning under uncertainty and highlighting its potential for broader application across diverse planning contexts in the energy sector.
Table 1: Main Notation
Symbol Meaning
$X=\{x_{1},\dots,x_{D}\}$ Investment decision variables
$\Xi$ High-dimmensional random variable over horizon $T$
$\hat{\Xi}$ Estimate of $\Xi$
$\xi$ Realization of uncertainty with respect to $\hat{\Xi}$
$Y$ Shortfall indicator
$G(x,\xi)$ System adequacy constraint
$C^{\text{inv}}(x)$ Investment cost
$C^{\text{op}}(x,\xi)$ Operational cost under $\xi$
$C^{\text{total}}(x)$ Total (investment + expected op.) cost
$p_{\Xi}(\xi)$ Density of stochastic variables
$p_{\Xi,Y,X}(\xi,y,x)$ Joint density of $(\Xi,Y,X)$
$p_{X}(x)$ Marginal density of $X$
$P_{\xi}(G(x,\xi)\geq 0)$ Chance constraint probability
$\varepsilon$ Adequacy tolerance
$\text{TPM}(\cdot)$ Tractable probabilistic model
$\mathcal{V},\mathcal{E}$ Nodes and edges of SPN graph
$\mathcal{V}_{\Sigma},\mathcal{V}_{\Pi},\mathcal{V}_{L}$ Sum, product, and leaf node sets
$O_{n}(x_{\psi(n)})$ Output of node $n$
$\psi(n)$ Scope (vars covered by node $n$)
$w_{a,n_{\Sigma}}$ Weight of child $a$ in sum node $n_{\Sigma}$
$ch(n)$ Set of children of node $n$
$o_{n}$ Log-probability output of node $n$
$m_{a,n_{\Sigma}}$ Binary var. for active child
$\text{T}^{\mathrm{LL}}_{n_{\Sigma}}$ Big-M constant for sum node
$K$ Number of scenarios
2 Background
Planning in Power Systems. Planning is a cornerstone of power system operation and development, ensuring that electricity supply can meet future demand reliably and at reasonable cost. Because power infrastructure has long lifetimes and requires significant capital investment, planning decisions must be made years or even decades in advance, often under deep uncertainty about technology, policy, and demand evolution. This task has traditionally been formulated as a capacity expansion problem, typically divided into generation expansion planning [DeJonghe2012830, Oderinwale2020] (GEP), transmission expansion planning [Xie2020, Lschenbrand2021] (TEP), and distribution expansion planning [Anjo2018, Gbadamosi2020] (DEP). Each has its own time horizons and decision variables. Collectively, these planning activities form the backbone of long-term power system development, linking near-term operational security with long-term goals such as decarbonization, resilience, and affordability [MOTTA2024401]. Traditional formulations often adopt deterministic structures that model uncertainty under a single problem, containing a set of scenarios, which are limited by computational power, leading to the common practice of selecting the most representative cases. As long-term planning increases uncertainty, it is typical to consider certain factors, such as focusing solely on one of the expansion problems, disregarding unit commitment constraints, or modeling the transmission system using a DC Optimal Power Flow (DCOPF) approach. The DCOPF is a linearized approximation of the AC power flow equations that simplifies the AC power flow equations, enabling scalable optimization for large planning problems, albeit at the expense of detailed network accuracy [doi:10.1049/iet-gtd.2015.0561]. More recent work emphasizes the need to explicitly capture uncertainty in demand and external drivers such as climate change. For example, the work in [LI2016401] models discrete climate scenarios within a stochastic generation expansion framework to evaluate viable investment strategies across multiple futures.
Reference Scenario Expansion. The formulation of the scenario expansion problem can have different forms depending on the specific case. Still, generally, it consists of first-stage (here-and-now) investment decisions and second-stage (wait-and-see) operational variables. For our reference case, let $X=\{x_{1},\dots,x_{D}\}$ denote the set of possible investment decisions, where $D$ is the number of available decisions. For instance, given a set of generation units $\mathcal{G}=\{g_{1},\dots,g_{G}\}$, a set of transmission lines $\mathcal{L}=\{\ell_{1},\dots,\ell_{L}\}$, and a set of storage units $\mathcal{S}=\{s_{1},\dots,s_{S}\}$, a single decision could have the form
$x_{d}=\Big\{$ $G,g_{1}^{\mathrm{cap}},\dots,g_{G}^{\mathrm{cap}},L,\ell_{1}^{\mathrm{cap}},\dots,\ell_{L}^{\mathrm{cap}},$
$S,s_{1}^{\mathrm{cap}},\dots,s_{S}^{\mathrm{cap}},\dots\Big\},$ (1)
establishing how many generation units to consider ($G$), and their capacity ($g_{1}^{\mathrm{cap}},\dots,g_{G}^{\mathrm{cap}}$), how many transmission lines would be needed ($L$), and their capacity ($\ell_{1}^{\mathrm{cap}},\dots,\ell_{L}^{\mathrm{cap}}$), and how many storage units ($S$) and their capacities ($s_{1}^{\mathrm{cap}},\dots,s_{S}^{\mathrm{cap}}$). Decisions are not limited to hardware; other potential components are portfolio or technology adoption choices, such as binary or continuous variables indicating whether a new technology is selected, or enforcing minimum/maximum shares of specific technologies.
The second group of variables represents short-term operational decisions, which tend to be related to stochastic factors such as meteorological events or demand, through a target time window of size $T$. We define $\Xi=\{\xi_{1}\oplus\dots\oplus\xi_{T}\}$ as concatenating all the stochastic variables after uncertainty is realized. A single instance for the timestep $t$ could be defined as follows:
$\xi_{t}=\Big\{$ $g^{\mathrm{out}}_{1,t},\dots,g^{\mathrm{out}}_{G,t},g^{\mathrm{curt}}_{1,t},\dots,g^{\mathrm{curt}}_{G,t},g^{\mathrm{res}}_{1,t},\dots,g^{\mathrm{res}}_{G,t},$
$\ell^{P}_{1,t},\dots,\ell^{\mathrm{P}}_{L,t},\ell^{Q}_{1,t},\dots,\ell^{\mathrm{Q}}_{L,t}s^{\mathrm{ch}}_{1,t},\dots,g^{\mathrm{ch}}_{S,t},$
$s^{\mathrm{dch}}_{1,t},\dots,g^{\mathrm{dch}}_{S,t},\dots\Big\},$ (2)
including variables like the actual generation output of the units ($g^{\mathrm{out}}_{1,t},\dots,g^{\mathrm{out}}_{G,t}$), curtailment ($g^{\mathrm{curt}}_{1,t},\dots,g^{\mathrm{curt}}_{G,t}$), reserve decisions ($g^{\mathrm{res}}_{1,t},\dots,g^{\mathrm{res}}_{G,t}$), operational values on the power lines such as active power ($\ell^{P}_{1,t},\dots,\ell^{\mathrm{P}}_{L,t}$) and reactive power ($\ell^{Q}_{1,t},\dots,\ell^{\mathrm{Q}}_{L,t}$), charging ($s^{\mathrm{ch}}_{1,t},\dots,g^{\mathrm{ch}}_{S,t}$) or discharging ($s^{\mathrm{dch}}_{1,t},\dots,g^{\mathrm{dch}}_{S,t}$) decisions on storage, among many other possibilities.
Together, $X$ and $\Xi$ incorporate uncertainty into the investment decision process to make a final decision that aligns with the operator’s primary objective: minimizing costs while maintaining the system’s reliability. However, as the space of possible values for each $\xi_{t}$ is extremely large, it is prohibitive to explore them all; thus, it is common to build an estimate ($\hat{\Xi}$) with some representative cases and optimize $x$ over it, yielding to solving one constrained optimization problem for each $\xi\sim\hat{\Xi}$ as follows:
$\min_{x}\quad$ $C^{\mathrm{inv}}(x)+C^{\mathrm{op}}(x,\xi)$ (3)
s.t. $G(x,\xi)\geq 0,$
$g(x)=0,$
$h(x)\leq 0,$
to posteriorly pick the least expensive solution among the selected scenarios. The constraints in (3) represent the various considerations that limit the minimization of the primary objective (investment cost) while ensuring compliance with all scenario-dependent operational and policy constraints. This compact notation allows the formulation to encompass various system constraints without explicitly enumerating each one.
Relevant Work on Scenario Expansion Planning. In power system planning, stochastic optimization and robust methods represent two complementary pathways to deal with uncertainty. For instance, Giannelos et al. [giannelos2025stochastic] propose a multi-stage stochastic network expansion model that accounts for both exogenous and endogenous uncertainties, capturing how investments evolve over time under varying conditions. On the robust side, Chen et al. [chenRobustOptimizationTransmission2014] introduced minimax cost and regret metrics to guard against worst-case outcomes. In contrast, Roldán et al. [roldanRobustTransmissionNetwork2019] extended that line by modelling correlated uncertainty sets, making the robust expansion plans more realistic. Other works further refine robustness: Li et al. [Li2016] incorporate ramping uncertainty and construction delays, and Velloso et al. [velloso2018distributionally] move toward distributionally robust formulations that sit between full stochastic and worst-case models. More recently, Brenner et al. [brenner2024learning] have combined stochastic planning with learning-based approximations to address computational complexity in large-scale systems. In parallel, Gholami et al. [gholami2022dynamic] explore joint expansion of power and gas infrastructures under uncertainty, illustrating how interdependent systems can be co-optimized.
Chance Constraint Optimization. A central difficulty in power system planning is that these decisions happen under uncertainty. Traditional deterministic formulations provided valuable insights but offered limited protection against the wide range of plausible futures. As the penetration of renewables grows and climate-related risks become more pronounced [paragElectricityMarketDesign2016a], planning frameworks increasingly require explicit methods to incorporate uncertainty and evaluate the robustness of proposed investment strategies. One promising direction is to include chance constraints, which involve enforcing constraints probabilistically rather than for all possible outcomes, an idea previously explored in operation optimization problems.
Chance-constrained formulations have already existed to enhance short-term scheduling. For example, Tan et al. [tan2020dual] developed a dual-timescale generation scheduling framework that integrates wind power, demand response, and energy storage, using chance constraints to procure sufficient ramping capacity under uncertainty. Similarly, Wang et al [wang2013stochastic] studied stochastic unit commitment with uncertain demand response, where a random variable models customer price elasticity and a chance constraint bounds the loss-of-load probability. For planning, Chen et al. [chen2024multistage] proposed a method for multi-stage planning that utilizes a chance constraint to address unmet demand.
Tractable Probabilistic Models. Probabilistic models have long provided a principled foundation for reasoning under uncertainty. These models describe a probability distribution over observed and latent variables, enabling one to answer questions such as: What is the likelihood of a given event occurring? What is the expected value of a function under uncertainty? Or what is the most likely explanation for the observed data? This ability to represent and manipulate uncertainty is essential for decision-making in domains ranging from natural language processing and computer vision to power systems and finance. In fact, probability theory remains the most widely adopted framework for decision-making under uncertainty in artificial intelligence (AI) and machine learning (ML).
Over the last two decades, the field has witnessed remarkable progress in developing expressive generative models. Deep generative approaches, such as generative adversarial networks [goodfellow2014gans], variational autoencoders [kingma2022autoencodingvariationalbayes], and normalizing flows [papamakarios2021normalizingflowsprobabilisticmodeling], enable closer approximation to real-world data density estimation. These models show outstanding performance at generating highly realistic samples and capturing intricate dependencies in data, offering unprecedented modeling flexibility. Additionally, probabilistic programming languages and advances in Bayesian nonparametrics have expanded the expressiveness of probabilistic modeling frameworks.
However, the increase in expressiveness comes at a high computational cost. Inference in these models (computing marginals, conditionals, expectations, or the most likely assignment of variables) is typically intractable. For most deep generative models, it is unfeasible to compute efficiently the probability of a given outcome; instead, inference relies on sampling or approximate optimization. These procedures can be slow, unstable, or provide no guarantees. As [choi2020probabilistic] highlights, this creates a paradox: while models become better at fitting data distributions with high fidelity, they become less reliable for reasoning about those distributions, undermining the very goal of probabilistic modeling, which is not only to mimic data but to support decision-making through accurate probability computations.
Tractable probabilistic models provide an alternative approach to addressing the aforementioned paradox. These models are designed explicitly to enable the exact computation of inference queries in polynomial time. The idea is not new: restricted forms of TPMs were proposed a few decades ago, such as the Kalman filter [Kalman1960], hidden Markov models (HMMs) [Baum1966], and tree-structured distributions [chow1968dependencetrees]. Although these models had limited expressive power, they were highly influential in signal processing, speech recognition, and computational biology thanks to their ability to support efficient inference algorithms.
Modern TPM research maintains the same philosophy: seeking greater expressiveness without sacrificing tractability. Some relevant methods include arithmetic circuits [darwiche2002logical], probabilistic decision diagrams [Jaeger2006, Mateescu2012], cutset networks [Rahman2014], and probabilistic sentential decision diagrams [kisa2014]. Among the most widely studied are sum–product networks (SPNs) [poon2011sum, xiaSurveySumProduct2023] and more general probabilistic circuits (PCs) [choi2020probabilistic, liu2021tpm], which represent distributions as layered computational graphs where sums encode mixtures and products encode independencies. These models provide a framework in which exact computations run in time proportional to the size of the circuit with some solid guarantees.
TPM vs. Other Probabilistic Models. Compared to mixture models (e.g., Gaussian mixtures), TPMs can scale to higher dimensions without exponential blow-up in the number of mixture components. Compared to deep generative models, TPMs offer sampling and exact probability evaluation, which is crucial in settings that require probabilistic guarantees—such as safety verification in autonomous systems, reliability analysis in power grids, or enforcing chance constraints in stochastic optimization. Monte Carlo simulation can approximate such probabilities, but it may require enormous sample sizes. In contrast, TPMs can compute them directly and precisely from the model representation.
TPMs in Chance-Constrained Optimization. TPMs allow efficient evaluation of probabilities, making them ideal candidates for implementing chance constraints in optimization. A notable recent example is LiCE [nemecek2024generating], which encodes an SPN into a mixed-integer optimization formulation, enabling constraints on the likelihoods of counterfactual explanations of AI models. Additionally, Ghandi et al. [ghandi2024probabilistic] propose methods to integrate propositional constraints directly into probabilistic circuits using convex optimization, refining the circuit distribution to satisfy logical constraints without retraining. Together, these works provide a foundation for embedding probabilistic constraints—such as enforcing reliability or safety with high probability—inside optimization models, suggesting TPMs could support chance-constrained expansion planning in energy systems.
3 Methodology
3.1 Chance-Constrained Capacity/Generation/Storage Expansion Problems
The traditional approach to solving a generation expansion problem, described in the previous section, relies on finding a global optimum of an approximated problem based on an estimate built from a set of representative realizations of the uncertainty $\hat{\Xi}$, in statistical terms, something like modeling the uncertainty as a mixture of Dirac distributions. Similar approaches in capacity expansion planning are utilized in [10931726]. The formulation in (3) requires finding the cheapest $x$ for which all sampled scenarios $\xi\sim\hat{\Xi}$ satisfy the system adequacy constraints simultaneously (or up to an adequacy level), leading to a computationally expensive solution when uncertainty is high. Probabilistic modeling is one alternative to the traditional approach, assuming access to a reliable model of $\mathbb{P}_{\xi\sim p(\Xi)}\left(G(x,\xi)\geq 0\right)$, which makes it possible to formulate a chance-constrained optimization problem, where constraint satisfaction is required only with a prescribed confidence level $(\varepsilon)$:
$\min_{x}\quad$ $C^{\mathrm{inv}}(x)+\mathbb{E}_{\xi\sim p(\Xi)}\left[C^{\mathrm{op}}(x,\xi)\right]$ (4)
s.t. $\mathbb{P}_{\xi\sim p(\Xi)}\left(G(x,\xi)\geq 0\right)\leq\varepsilon,$
$g(x)=0,$
$h(x)\leq 0.$
Chance-Constrained Capacity/Generation/Storage Expansion Problems with TPMs. The main challenges that emerge from the problem described in (4) are two-fold: 1) we need a reliable model for the chance constraint, and 2) we need to be able to perform inference over it many times. Fortunately, TPMs offer a feasible alternative to tackle both challenges, first, by allowing for the efficient estimation of the chance constraint, and second, by enabling integration into optimization pipelines, as shown in the work of [nemecek2024generating]. To enable the integration in such a framework, we start by defining a binary random variable $Y=\mathds{1}(G(x,\xi)\geq 0)$ which allows expressing the chance constraint as follows:
$\mathbb{P}_{\xi\sim p(\Xi)}\left(G(x,\xi)\geq 0\right)$ $=\mathbb{P}_{\xi\sim p(\Xi),x\sim X}\left(G(x,\xi)\geq 0|X=x\right)$
$=\int_{-\infty}^{\infty}p_{\Xi\mid X}(\xi|x)\mathds{1}(G(x,\xi)\geq 0)d\xi$
$=\int_{-\infty}^{\infty}p_{\Xi,Y\mid X}(\xi,1|x)d\xi$
$=\int_{-\infty}^{\infty}\frac{p_{\Xi,Y,X}(\xi,1,x)}{p_{X}(x)}d\xi.$ (5)
The reformulation (5) shows that the chance constraint could be estimated by approximating joint and marginal probability densities over the random variables $X$, $\Xi$, and $Y$. As mentioned before, TPMs allow inference queries to approximate densities in polynomial time, enabling us to compute the ones we need to estimate the chance constraint just by using different queries to the TPM:
$\text{TPM}(\xi,y,x)$ $\approx p_{\Xi,Y,X}(\xi,y,x),$ (6)
$\text{TPM}(y,x)$ $\approx p_{Y,X}(y,x),$ (7)
$\text{TPM}(x)$ $\approx p_{X}(x),$ (8)
where, depending on the query, one can compute an estimate for the joint distribution of all the random variables (6), marginalize over a variable, for instance over $\Xi$ in (7), or compute the marginal of one of the variables, such as in (8) for $X$. Using a TPM, we can reformulate the chance constraint as
$\frac{\text{TPM}(x,y=1)}{\text{TPM}(x)}\leq\varepsilon.$ (9)
Another way of understanding (9) is that, using an arbitrary number of samples $(x,\xi)$, we train a TPM to learn a joint distribution over the stochasticity. With that joint distribution, we marginalize and condition over given evidence to compute different densities. Furthermore, if we can sample $x$ from a given distribution, we can replace the denominator in (9) by $p_{X}(x)$ directly.
Optimizing TPM-Based Chance Constraint. In the energy industry, most large-scale planning and operational problems are formulated as linear or mixed-integer linear programs (LP/MILP) because of their computational efficiency and compatibility with commercial solvers. These formulations dominate applications such as generation and transmission expansion, unit commitment, and market clearing [conejo2010decision, kirschen2018fundamentals]. Nonlinear formulations—such as AC optimal power flow—are typically linearized to maintain convexity and guarantee scalability. More recent work has extended these methods through stochastic and robust optimization frameworks [chenRobustOptimizationTransmission2014, roldanRobustTransmissionNetwork2019], enabling planners to explicitly account for uncertainty while preserving the tractable structure required for practical implementation. It is natural to require that our proposed chance-constrained approach be compatible with LP/MILP solvers. As shown in the work of Němecěk et al. [nemecek2024generating], we can achieve this by selecting a TMP that can be translated into linear constraints, in their case, an SPN, which is our chosen approach as well. An SPN is a structured probabilistic model that represents a complex joint distribution through a combination of weighted sums and products of simpler distributions.
Formally, an SPN is a type of probabilistic circuit defined by a tuple $(\mathcal{G},\psi,\theta)$, where $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a directed acyclic graph (DAG) describing the computation structure ($\mathcal{V}$ is a set of vertices (nodes) and $\mathcal{E}$ the edges that interconnect them), $\psi:\mathcal{V}\rightarrow 2^{[P]}$ is a scope function that assigns to each node a subset of input features which define its distribution (i.e. the root node $\mathcal{V}^{\mathrm{root}}$ covers the entire set of random variables), and $\theta$ denotes the model parameters. For each inner node $n\in\mathcal{V}$, we define $\mathrm{ch}(n)$ as the set of its child nodes and denote by $x_{\psi(n)}$ the input features that belong to the scope of node $n$.
A leaf node $n^{L}\in\mathcal{V}^{L}=\{n\mid\mathrm{ch}(n)=\emptyset\}$ represents a probability distribution over its local scope,
$O_{n^{L}}(x_{\psi(n^{L})})=p(x_{\psi(n^{L})};\theta_{n^{L}}),$ (10)
where $p(\cdot)$ is a tractable distribution parameterized by $\theta_{n^{L}}$. A product node $n^{\Pi}\in\mathcal{V}^{\Pi}$ performs a product over its children,
$O_{n^{\Pi}}(x_{\psi(n^{\Pi})})=\prod_{a\in\mathrm{ch}(n^{\Pi})}O_{a}(x_{\psi(a)}).$ (11)
To ensure tractability, the scopes of the children of a product node must be decomposable, meaning that they are disjoint and collectively cover the parent’s scope:
$\bigcup_{a\in\mathrm{ch}(n^{\Pi})}\psi(a)=\psi(n^{\Pi}),\qquad\bigcap_{a\in\mathrm{ch}(n^{\Pi})}\psi(a)=\emptyset.$ (12)
A sum node $n^{\Sigma}\in\mathcal{V}^{\Sigma}$ computes a weighted mixture of the distributions defined by its children:
$O_{n^{\Sigma}}(x_{\psi(n^{\Sigma})})=\sum_{a\in\mathrm{ch}(n^{\Sigma})}w_{a,n^{\Sigma}}\,O_{a}(x_{\psi(a)}),$ (13)
where each weight must be non-negative ($w_{a,n^{\Sigma}}\geq 0$) and $\sum_{a\in\mathrm{ch}(n^{\Sigma})}w_{a,n^{\Sigma}}=1$. Additionally, all children of a sum node must have identical scopes (a property known as completeness or smoothness):
$\psi(a_{1})=\psi(a_{2}),\qquad\forall\,a_{1},a_{2}\in\mathrm{ch}(n^{\Sigma}).$ (14)
These structural constraints for sum nodes ensure that SPNs support exact marginalization and conditioning in linear time with respect to the network size. As a result, SPNs can efficiently compute densities and expectations required in stochastic and chance-constrained optimization problems. However, to enable the use of an SPN in the LP/MILP setup, we need to encode them into linear constraints, which is non-trivial, and some additional considerations are required. Following the approach of Němeček et al. [nemecek2024generating], changing all computations in logarithmic space simplifies the formulation of product nodes and improves numerical stability.
Starting with leaf nodes $n^{L}\in\mathcal{V}^{L}$, they represent a tractable univariate probability distribution. For discrete variables, one can use indicator variables $d_{j,k}$ to encode the active category $k$ of feature $j$ in a leaf node. For continuous variables, we can use a piecewise-linear approximation (e.g., histograms or linearized density functions) to represent the log-likelihood within a linear model. These leaf evaluations are denoted by $o_{n^{L}}$, representing the log-probability output of node $n^{L}$.
For the case product node $n^{\Pi}\in\mathcal{V}^{\Pi}$ aggregates the log-probabilities of its children additively. Let $\mathrm{ch}(n^{\Pi})$ denote the set of predecessors of node $n^{\Pi}$, each providing an output $o_{a}$. The output of a product node in log-space is then:
$o_{n^{\Pi}}=\sum_{a\in\mathrm{ch}(n^{\Pi})}o_{a},\qquad\forall n^{\Pi}\in\mathcal{V}^{\Pi}.$ (15)
This relationship is purely linear and can be directly applied to linear programming.
The sum nodes $n^{\Sigma}\in\mathcal{V}^{\Sigma}$ compute a weighted mixture of their children’s outputs. In probability space, this corresponds to a convex combination $\sum_{a\in\mathrm{ch}(n^{\Sigma})}w_{a,n^{\Sigma}}\,\exp(o_{a})$; this term is nonlinear in log-space. To retain tractability, we approximate the logarithm of the sum by its dominant term using a “log-sum-exp” upper bound, which can be expressed as
$o_{n^{\Sigma}}\approx\max_{a\in\mathrm{ch}(n^{\Sigma})}\left(o_{a}+\log w_{a,n^{\Sigma}}\right).$ (16)
This approximation can be linearized using binary variables $m_{a,n^{\Sigma}}\in\{0,1\}$ to indicate which child is active. We define $\forall a\in\mathrm{ch}(n^{\Sigma}),\;n^{\Sigma}\in\mathcal{V}^{\Sigma}$ the following:
$o_{n^{\Sigma}}$ $\leq o_{a}+\log w_{a,n^{\Sigma}}+m_{a,n^{\Sigma}}\,\text{T}^{\mathrm{LL}}_{n^{\Sigma}},$ (17)
where $T^{\mathrm{LL}}_{n^{\Sigma}}$ is a large constant (“big-M”) ensuring that only one child can determine the active maximum, achieved through the following additional constraint:
$\sum_{a\in\mathrm{ch}(n^{\Sigma})}m_{a,n^{\Sigma}}$ $=|\mathrm{ch}(n^{\Sigma})|-1,$ $\forall n^{\Sigma}\in\mathcal{V}^{\Sigma}.$ (18)
In this formulation, the approximation provides a lower bound on the actual log-sum-exp value, which is sufficient when maximizing likelihood or probability mass.
Overall, these constraints encode the structure of the SPN within a mixed-integer linear program. Product nodes remain linear, while sum nodes require a small number of binary variables per mixture component. This formulation enables the direct inclusion of probabilistic reasoning inside optimization models, allowing the evaluation of likelihood-based or chance-constrained objectives using standard MILP solvers. Table 1 summarizes the most relevant notation used in this section.
Figure 1: Illustrative comparison of computational scaling between traditional scenario-based formulations and the proposed SPN-based chance-constrained approach. While scenario-based methods scale linearly or superlinearly with the number of scenarios, the SPN-based formulation requires only a one-time training cost and remains effectively constant afterward.
Computational Scaling and Scenario Independence. Its favorable computational scaling is a key advantage of reformulating the problem as a chance-constrained optimization with a tractable probabilistic model. In traditional scenario-based formulations, the number of variables and constraints grows proportionally to the number of sampled uncertainty realizations, i.e., $\mathcal{O}(K)$. This linear (or sometimes superlinear) growth rapidly increases the computational burden as the number of scenarios $K$ grows, since each additional scenario introduces a new set of operational constraints. As illustrated in Figure 1, this significantly increases solver time, especially when thousands of scenarios are required to capture high-dimensional uncertainty.
In contrast, the proposed TPM-based formulation decouples the optimization problem from the explicit enumeration of scenarios. With a trained SPN, evaluating or marginalizing probabilities required for the chance constraint involves only a fixed number of inference queries, each computable in time linear with respect to the network size but independent of $K$ and effectively replacing scenario enumeration with probabilistic inference, enabling the optimizer to reason over the full uncertainty distribution without increasing the model’s dimensionality. Moreover, because TPMs train on arbitrary numbers of samples, their accuracy improves with larger datasets; thus, unlike scenario-based methods, increasing $K$ enhances model quality rather than incurring additional computational cost.
Figure 2: A schema of the proposed framework.
3.2 TPM-Based Pipeline for Probabilistic Energy System Planning
To demonstrate that TPMs can solve the Capacity/Generation/Storage Expansion problems, we propose a pipeline organized into three layers as shown in Figure 2. The first layer is a realistic simulator representing a generation system design scenario, in which decisions (based on hardware specifications) are made in conjunction with prototypical household load shapes and stochastic environmental variables, determining whether the system will fulfill its demands successfully or suffer shortfalls. The simulator is a scenario generator that produces large numbers of daily realizations. The second layer performs probabilistic learning using SPNs, which capture the conditional probability of shortfall events for each possible decision. The third layer couples these probabilistic outputs with an MILP engine, enabling risk-aware design by imposing chance constraints on shortfall probability.
3.2.1 Layer 1: Generation Design Simulator
The first layer represents a simulator of the power generation and storage system, formalized as a mapping:
$f:(x,\xi)\;\mapsto\;y,$ (19)
where $x\sim X$ denotes the vector of decision variables (e.g., number of PV panels, number of battery units, inverter type, and control policy), $\xi\sim\Xi$ the realization of the stochastic variables (e.g., irradiance, temperature, and load profile), and $y\sim Y$ the adequacy outcome indicating whether the system meets demand ($y=0$) or experiences a shortfall ($y=1$).
Given a configuration $x$ and a realization of the stochastic environment $\xi$, the simulator computes the system’s power balance across a planning horizon, accounting for generation, storage operation, and load demand. We label a scenario as a shortfall when the instantaneous deficit between generation and demand exceeds a predefined tolerance. By repeatedly evaluating $f(x,\xi)$ under many draws of $\xi$, the simulator generates a large dataset of labeled daily realizations:
$\mathcal{D}=\{(x_{i},\xi_{i},y_{i})\}_{i=1}^{N},$
which serves as the empirical distribution from which we train the tractable probabilistic model in Layer 2.
This representation captures how investment and control decisions interact with environmental and demand uncertainty to determine system adequacy while abstracting away the internal details of physical modeling. The resulting dataset provides a statistically rich foundation for probabilistic learning, ensuring that the optimization layer can reason about reliability without explicit scenario enumeration.
3.2.2 Layer 2: SPN Training and Inference
The second layer learns a tractable probabilistic representation of the relationship between design choices, environmental factors, and adequacy constraints using the labeled dataset generated in Layer 1. We train an SPN to approximate the joint density $p(X,\Xi,Y)$, where $X$ represents design and control variables, $\Xi$ the stochastic environmental inputs, and $Y$ the adequacy label (shortfall or no-shortfall). Once trained, the SPN supports exact density queries such as the conditional probability of system adequacy given a specific design, $\mathbb{P}(Y=1\mid X=x)$, and marginal probabilities across subsets of variables. Because SPNs permit inference in polynomial time with respect to the network size, the resulting model enables the evaluation of chance constraints over arbitrarily large scenario spaces without explicit enumeration. This layer bridges the data-driven simulation environment with the optimization stage by providing an interpretable and tractable probabilistic surrogate for system reliability.
3.2.3 Layer 3: MILP Integration and Optimization
The third layer integrates the trained SPN into an MILP framework to perform risk-aware capacity, generation, and storage planning. Notice that in the chance constraint (5), one could consider the probability of “inadequacy”, i.e., the event where the demand exceeds generation by more than some tolerance $\varepsilon$. The positive part of demand with the generation subtracted is sometimes also known as the instantaneous shortfall. The optimization seeks a configuration $x$ that minimizes total investment and operational costs while ensuring that the probability of inadequacy remains below a predefined threshold throughout the planning horizon. For practical purposes, in our methodology, we will reduce the objective to only the investment component:
$\min_{x}\quad$ $C^{\mathrm{inv}}(x)$ (20)
s.t. $\mathbb{P}_{\xi\sim p(\Xi)}(\text{instantaneous shortfall}\geq\text{tolerance})<\varepsilon.$
The SPN evaluates the probability term directly, whose log-space formulation can be encoded as a set of linear and mixed-integer constraints, as detailed in Section 3.1. This encoding allows the optimizer to compute feasible and near-optimal solutions using standard MILP solvers without resorting to Monte Carlo sampling. By coupling tractable probabilistic inference with deterministic optimization, this layer enables the selection of system designs that minimize cost while guaranteeing shortfall probabilities within acceptable limits.
4 Experimetal Results
4.1 Probability Landscape Comparison
To evaluate the ability of SPN-based methods to learn the true probability distribution over system configurations, we compare their probability landscapes with the empirical estimation of our probability of interest (20) across different levels of data availability. We aimed to verify whether high-probability regions, ridges, and exclusion zones are accurately recovered in the correct locations, how estimates evolve with additional data (convergence), and where systematic biases or overconfidence may persist.
Figure 3: Comparison between empirical and SPN-estimated shortfall ratios (SPN, SPN-Max, and SPN-Piecewise): In order, the Empirical column shows the best empirical estimation possible with respect to access to more samples. The SPN column displays the learned approximation. The SPN-Max column shows the first proposed approximation (SPN-Max). The SPN-Piecewise column displays the improved version of our proposed approximation. We highlight that our best method is capable of capturing the pattern of the best empirical approximation. Although SPN-Max captures the pattern as well, the error scales with the size of the SPN, something that is minimized with SPN-Piecewise while remaining computable using MILP.
The key analytical value of the grid layout in Figure 3 lies in enabling two critical comparisons simultaneously: horizontal comparison within each row reveals whether the SPN-based methods accurately capture the probability patterns of the best empirical estimator for a given data availability scenario, while vertical comparison across rows shows how these probability estimates evolve and potentially converge toward ground truth as more training data becomes available. Ideally, the heatmaps in each row should exhibit similar color patterns, with bright yellow regions (indicating high probability) appearing in the same locations across all methods, suggesting that the SPN has correctly learned which configurations correspond to a low likelihood of a shortfall event.
The results show that SPN progressively aligns more closely with the empirical reference patterns, reflecting the improvement in the model’s ability to represent the desired density as the training data increases. The comparison between SPN (expectation over all paths) and SPN-Max (maximum probability path) reveals important characteristics about the network structure: significant differences between these two suggest that the probability mass is distributed across multiple paths, while near-identical patterns indicate the network relies heavily on a dominant path, demonstrating that the approximation in SPN-Max comes at the cost of imprecision in cases where there is not a dominant path in sum nodes. As a consequence, the error grows with the number of nodes of the SPN. The SPN-Piecewise approximation significantly reduces the error. We can conclude that poor optimization decisions stem from inaccurate probability modeling (misaligned heatmap patterns) or from issues in the optimization procedure itself (correct probability patterns but suboptimal configuration selection).
Figure 4: Configuration space exploration across data sparsity levels (30 iterations per percentage). Point clouds show each method’s configuration choices with capacity-based axes (primary) and unit counts (secondary). A gold star indicates the ground truth optimum. Red dashed boundary marks the region where the adequacy constraint ($\mathbb{P}(\text{shortfall})=\varepsilon$, where $\varepsilon=5\%$) is respected based on ground truth. Configurations outside this boundary fail the adequacy requirement. Note that these methods may appear safe based on their own estimates, yet select configurations that exceed the true safety boundary, revealing calibration overconfidence. We added some jitter to the points to show the clusters of runs, but they concentrate around the same solution.
4.2 Reliability of SPN-Based Investment Planning under Data Availability Constraints
In this experiment, we analyze how the reliability of SPN–based investment planning evolves as a function of the amount of training data available, following the insight in section 4.1. The goal is to evaluate whether the probabilistic representation learned by the SPN can accurately capture the operational boundaries of the system, with a primary objective of respecting the adequacy constraint, which limits the probability of experiencing power shortfalls below a safety threshold (e.g., $\mathbb{P}(\text{shortfall})\leq\varepsilon$, where $\varepsilon=5\%$).
We conducted simulations across five dataset availability levels, corresponding to $10\%$, $25\%$, $50\%$, $75\%$, and $100\%$ of the total generated dataset (45,000 realizations). For each level, and for 30 iterations to ensure statistical robustness, we executed our pipeline: 1) randomly sampling a smaller version of the dataset, 2) fitting an SPN using the best hyperparameters we found through a Bayesian search, and 3) obtaining the optimal solution that minimizes (20). The experiment’s design proposes a realistic planning scenario in which data availability is limited, for instance, when only a subset of operational histories or synthetic realizations is accessible for training. We kept as a reference the complete dataset, which yields the best estimations we could access of our adequacy constraint. Finally, we compared four different optimization methods:
  • GEP10: Following the notation (3), we built an estimation ($\hat{\Xi}$) from $K=10$ representative scenarios, ran them by our simulator, and optimized a problem as follows:
    $\min_{x}\quad$ $C^{\mathrm{inv}}(x)$ (21)
    s.t. $G(x,\xi)\geq 0,\quad\text{with}~\xi\sim\hat{\Xi}$
    We assumed that optimizing a larger set is prohibitive, and the objective aims to minimize the investment in a configuration that meets the adequacy requirement. This formulation tends to underestimate risk because of its limited stochastic coverage.
  • GEP50: Similar to GEP10, but with $K=50$ representative scenarios. Increasing the number of scenarios improves the stochastic representation, typically producing safer (though potentially more expensive) solutions.
  • SPN: Uses the fitted SPN (using the available data sampled for the specific iteration) to directly estimate the probability of adequacy violations. This estimation cannot be the input of an optimizer, but it is still a point of reference for what the best solution the SPN-based strategies could yield.
  • SPN-Max: This value follows the estimation proposed in [nemecek2024generating] to have an MILP-compatible function. This approach enables the embedding of SPNs within a tractable optimization pipeline, allowing for the explicit enforcement of chance constraints during decision-making.
  • SPN-Piecewise: This is an improved formulation of SPN using MILP constraints, using piece-wise linear functions to approximate the exponential and logarithm functions. For each sum node $n^{\Sigma}$, the maximal input value is subtracted from all values for numerical stability; then, we approximate the exponentials on the range $\ln 0.001$, 0] where values are numerically relevant (values below 0.001 are rounded to 0). Then, a logarithm approximation of the sum of exponential outputs is taken, where the logarithm is approximated on [1, $|\mathrm{ch}(n^{\Sigma})|$] to cover all possible values. Each function is approximated using five linear segments. The output of the logarithm is finally summed with the subtracted maximal value, providing a tighter approximation of the true value.
Each subplot in Figure 4 corresponds to one of these dataset fractions. The horizontal axis represents the battery capacity (in MWh), while the vertical axis corresponds to the solar generation capacity (in MW). The red dashed contour marks the safety boundary, which separates configurations that meet the adequacy constraint from those that do not. The star denotes the ground truth optimal configuration, obtained from the full-fidelity simulation model that explicitly enumerates all stochastic outcomes.
This experiment yields important insights that validate our method, while also highlighting key considerations for future research. First, it illustrates the progressive improvement of the SPN’s reliability with larger dataset sizes. A limited number of samples allows fitting an SPN roughly; it will learn some high-level patterns that might allow finding a feasible region to make a decision, but as it receives more data, it will undergo a process of refinement that could damage the quality of the estimation (like in the case of $50\%$). The SPN-based approximations depend directly on how well the SPN is fitted, which will be posteriorly encoded in the MILP. However, our test showed that the error introduced by the approximation of the sum node might have a regularization effect (causing the decision to be conservative), which, as a consequence, keeps the decision consistently within the allowance constraint throughout the different data availability scenarios, which, in our experiment, is essential.
Comparing the SPN approach with GEP10 and GEP50 reveals several advantages to using SPNs. In our experimental setup, traditional GEP approaches heavily depend on the quality of the selected scenarios, which we test by selecting them from the available samples at each iteration of our experiment. When choosing a few scenarios, there is more uncertainty in fulfilling the adequacy constraint, as evidenced in Figure 4, where GEP10 shows large variability among the iterations and provides no guarantees of preventing shortfall. Switching to GEP50, considering a larger set of reference scenarios improves the consistency of the decision; however, it does not offer certainty about respecting the adequacy either. In both cases, the quality of the available scenarios becomes a matter.
We noticed that GEP10 exhibits high variability in the cost of the selected solution, which is a consequence of its dependence on scenario selection and the limitation on the number of scenarios that can be considered ($K$). The case of GEP50 shows an expected outcome: solutions would tend to be more expensive than the optimal one, but they show the lowest deviation from the optimal cost. However, the average solution lies on the boundary of adequacy, meaning that any error in the selection or simulation could cause the solution to move outside the safe area. Finally, the SPN methods tend to be conservative in their decisions, a result of the quality of the fitting process and the error introduced by the approximation. SPN methods demonstrate greater consistency across various data availability scenarios and random samples. This is a clear advantage, considering that the operation cost component $C^{\mathrm{op}}(x,\xi)$ in (4) would be negatively affected by shortfall events. Thus, we argue that TPM-based methods for solving the chance-constrained Capacity/Generation/Storage expansion problem offer a safer and more cost-effective (in the long term) alternative to traditional methods.
4.3 Training Details
We trained our SPNs using the SPFlow library [Molina2019SPFlow], which supports automatic structure learning and exact inference for tractable probabilistic models through queries. The structure learning procedure relies on hierarchical clustering and recursive variable partitioning to identify conditional independencies within the data. We optimized the hyperparameters using WandB [wandb] through Bayesian search over 500 trials, per percentage of the total dataset, with model selection based on ground-truth log-likelihood on the validation set. We standardized and discretized features depending on the leaf node type. Individual runs required from a few seconds to several minutes, depending on the dataset size and model complexity. Early stopping based on validation likelihood was employed to prevent overfitting. To ensure full reproducibility, we plan to make all code, configuration files, and instructions publicly available, thereby facilitating the replication of the experiments.
5 Conclusion
This work presents a probabilistic framework for investment planning in power systems under uncertainty, based on Sum–Product Networks (SPNs), by embedding these SPNs within a mixed-integer linear programming (MILP) formulation. Based on existing work on Tractable Probabilistic Models (TPMs), we demonstrated that probabilistic reasoning can be incorporated into chance-constrained optimization problems without explicit scenario enumeration, thereby redefining the traditional capacity/generation/storage scenario expansion problem by integrating a probabilistic adequacy constraint.
Our experimental results on a representative generation–storage planning case study indicated that SPN-based models can accurately capture adequacy probabilities and maintain reliable performance even under limited data availability. Compared to traditional scenario-based methods, the SPN formulation yielded more consistent and conservative investment decisions, mitigating shortfall risk at moderate computational cost. These findings highlight the potential of TPMs as a foundation for scalable, data-driven, and reliability-aware expansion planning in future low-carbon power systems that rely on stochastic, non-dispatchable sources. This work shows that TPMs offer a promising and interpretable approach to integrating uncertainty quantification and optimization in long-term power system planning.
6 Limitations
While the proposed framework exhibits strong potential, several significant limitations are worth highlighting, which could facilitate further research. First, the SPN linearization introduces an approximation that may lead to conservative or locally imprecise density estimation, yielding a calibration problem and potentially introducing overconfidence when multiple probabilistic paths contribute significantly to the overall likelihood. Future work will explore alternative formulations, other probabilistic circuits, or hybrid approaches to reduce this approximation error. Additionally, the quality of probabilistic inference depends on the representativeness of the training data. Real-world datasets may exhibit bias or insufficient coverage of rare events, which could affect reliability estimation. Incorporating active learning or uncertainty-aware sampling could improve robustness in such settings. Although the TPM formulation scales more favorably than scenario-based methods, embedding large SPNs within MILP solvers can pose computational challenges for full-scale systems if the learned structure becomes increasingly complex.